\begin{thebibliography}{99}




\bibitem{Bekenstein}
J. D. Bekenstein, Transcendence of the law of baryon-number conservation in black hole physics, Phys. Rev. Lett. 28, 452 (1972).




\bibitem{Chase}
J. E. Chase, Event horizons in Static Scalar-Vacuum Space-Times, Commun. Math. Phys. 19, 276 (1970).



\bibitem{C. Teitelboim}
C. Teitelboim, Nonmeasurability of the baryon number of a black-hole, Lett. Nuovo Cimento 3, 326 (1972).





\bibitem{Ruffini-1}
R. Ruffini and J. A. Wheeler, Introducing the black hole, Phys. Today 24, 30 (1971).




\bibitem{Hod-2}
S. Hod,Stationary resonances of rapidly-rotating Kerr black holes, The Euro. Phys. Journal C 73, 2378 (2013).




\bibitem{Hod-5}
S.Hod, The superradiant instability regime of the spinning Kerr black hole, Phys. Lett. B 758, 181(2016).


\bibitem{c1}
Carlos Herdeiro,Vanush Paturyan,Eugen Radu, D.H. Tchrakian,Reissner-Nordstr$\ddot{o}$m black holes with non-Abelian hair,
Phys. Lett. B 772(2017)63-69.

\bibitem{c2}
Maur\'{\i}cio Richartz, Carlos A. R. Herdeiro, Emanuele Berti,
Synchronous frequencies of extremal Kerr black holes: resonances, scattering and stability,Phys. Rev. D 96, 044034 (2017).

\bibitem{c3}
Yves Brihaye, Carlos Herdeiro, Eugen Radu, D.H. Tchrakian,
Skyrmions, Skyrme stars and black holes with Skyrme hair in five spacetime dimension,JHEP 1711(2017)037.




\bibitem{Benone}
C. L. Benone, L. C. B. Crispino, C. Herdeiro, and E. Radu, Kerr-Newman scalar clouds, Phys.
Rev. D 90, 104024 (2014).




\bibitem{Herdeiro-2}
C. Herdeiro, E. Radu, and H. R\'{u}narsson, Non-linear QQ-clouds around Kerr black holes, Phys. Lett. B 739, 302 (2014).






\bibitem{mr3}
Elizabeth Winstanley,Classical Yang-Mills black hole hair in anti-de Sitter space, Lect.Notes Phys.769:49-87,2009.



\bibitem{NSHB}
Yan Peng, Hair mass bound in the black hole with non-zero cosmological constants,
Physical Review D 98,104041,(2018).




\bibitem{NSHB1}
Yan Peng,Hair distributions in noncommutative Einstein-Born-Infeld black holes,Nucl. Phys. B 941(2019)1-10.







\bibitem{Yves Brihaye}
Yves Brihaye, Thomas Delplace, Carlos Herdeiro, Eugen Radu,
An analytic effective model for hairy black holes,
Phys. Lett. B 782(2018)124-130.



\bibitem{Rogatko2}
Marek Rogatko,Uniqueness of higher-dimensional phantom field wormholes,
Phys. Rev. D 97(2018)no.2,024001.





\bibitem{Degollado}
J. C. Degollado and C. A. R. Herdeiro, Stationary scalar configurations around extremal charged black holes,
Gen. Rel. Grav. 45, 2483 (2013).


\bibitem{Cunha}
P. V. P. Cunha, C. A. R. Herdeiro, E. Radu, and H. F. R\'{u}narsson, Shadows of Kerr black holes with scalar hair, Phys.
Rev. Lett. 115, 211102(2015).






\bibitem{Brihaye}
Y. Brihaye, C. Herdeiro, and E. Radu, Inside black holes with synchronized hair, Phys. Lett. B 760, 279 (2016).


\bibitem{Bekenstein-1}
J. D. Bekenstein, Black hole hair: 25-years after, arXiv:gr-qc/9605059.


\bibitem{CAR}
Carlos A. R. Herdeiro, Eugen Radu,Asymptotically flat black holes with scalar hair: a review, Int.J.Mod.Phys.D
24(2015)09,1542014.

\bibitem{Hod-6}
S.Hod, No-scalar-hair theorem for spherically symmetric reflecting stars, Physical Review D 94, 104073 (2016).




\bibitem{Hod-7}
S.Hod, No nonminimally coupled massless scalar hair for spherically symmetric neutral reflecting stars,Physical Review D 96, 024019 (2017).



\bibitem{Bhattacharjee}
Srijit Bhattacharjee, Sudipta Sarkar, No-hair theorems for a static and stationary reflecting star, Physical Review D 95, 084027 (2017).









\bibitem{Hod-8}
S.Hod, Charged massive scalar field configurations supported by a spherically symmetric
charged reflecting shell,Physics Letters B 763, 275 (2016).



\bibitem{Hod-9}
S.Hod, Marginally bound resonances of charged massive scalar fields in the background of
a charged reflecting shell, Physics Letters B 768(2017)97-102.


\bibitem{Yan Peng-1}
Yan Peng, Bin Wang, Yunqi Liu, Scalar field condensation behaviors around reflecting shells in
Anti-de Sitter spacetimes, Eur.Phys.J. C 78 (2018) no.8, 680.



\bibitem{bm-1}
S. Hod, Stability of the extremal Reissner-Nordstr$\ddot{o}$m black hole to charged scalar perturbations, Phys. Lett. B
713, 505 (2012).

\bibitem{bm-2}
S. Hod,No-bomb theorem for charged Reissner-Nordstr$\ddot{o}$m black holes, Phys. Lett. B 718, 1489 (2013).




\bibitem{Hod-10}
Shahar Hod,Charged reflecting stars supporting charged massive scalar field configurations, European Physical
Journal C 78, 173 (2017).


\bibitem{Yan Peng-2}
Yan Peng, Scalar field configurations supported by charged compact reflecting stars in a curved spacetime,
Physics Letters B 780(2018)144-148.



\bibitem{Yan Peng-3}
Yan Peng, Scalar condensation behaviors around regular Neumann reflecting stars, Nucl.Phys. B 934 (2018) 459-465.




\bibitem{Yan Peng-4}
Yan Peng, Static scalar field condensation in regular asymptotically AdS reflecting star backgrounds,
Phys. Lett. B 782(2018)717-722.



	
\bibitem{Yan Peng-5}
Yan Peng, On instabilities of scalar hairy regular compact reflecting stars, JHEP 10(2018)185.


\bibitem{Yan Peng-6}
Yan Peng, Hair formation in the background of noncommutative reflecting stars, Nucl.Phys. B 938(2019)143-153.




\bibitem{YanD}
Yan Peng, No hair theorem for spherically symmetric regular compact stars with Dirichlet boundary conditions,
Phys. Lett. B 792(2019)156-159.



\bibitem{LR}
Yan Peng, Large regular reflecting stars have no scalar field hair, Eur. Phys. J. C 79(2019)no.4,309.




\bibitem{LR1}
Yan Peng, No scalar hair theorem for neutral Neumann stars: static massive scalar fields nonminimally coupled to gravity,
arXiv:1903.01713[gr-qc].











\bibitem{FE}
F.E. Schunck, E.W. Mielke, General relativistic boson stars, Class. Quant. Grav. 20(2003)R301-R356.











\bibitem{DAER}
Dumitru Astefanesei, Eugen Radu, Boson stars with negative cosmological constant,
Nucl. Phys. B 665(2003)594-622.






\bibitem{GHD}
G.H. Derrick, Comments on nonlinear wave equations as models for elementary particles, J. Math. Phys. 5(1964)1252-1254.





\bibitem{ng1}
S.Hod, No-go theorem for static boson stars, Physics Letters B 778(2018)239-241.





\bibitem{ng2}
S. Hod, No-go theorem for spatially regular boson stars made of static
nonminimally coupled massive scalar fields, Eur. Phys. J. C 79(2019)no.1,26.




\bibitem{ng3}
Yan Peng, No-go theorem for static boson stars with negative cosmological constants,
Nucl.Phys.B 953(2020)114955.


\bibitem{cond1}
Filipp N. Rybakov, Egor Babaev, The absence of superconductivity in the next-to-leading order Ginzburg-Landau
functional for Bardeen-Cooper-Schrieffer superconductor, arXiv:2106.0263.





\bibitem{cond2}
A. V. Vagov, A. A. Shanenko, M. V. Milo\v{s}evi\'{c},
V. M. Axt, and F. M. Peeters, Extended ginzburg-landau formalism: Systematic expansion in
small deviation from the critical temperature,
Phys.Rev.B 85,014502(2012).





\bibitem{cond3}
A. Vagov, A. A. Shanenko, M. V. Milo\v{s}evi\'{c}, V. M. Axt, V. M. Vinokur, J. A. Aguiar,
and F. M. Peeters, Superconductivity between standard types: Multiband versus
single-band materials, Phys. Rev. B 93,174503(2016).








\bibitem{GLYP}
Guohua Liu, Yan Peng, No scalar condensations outside reflecting stars with coupling terms from Ginzburg-Landau models,
arXiv:2203.15564 [gr-qc].




\bibitem{dg}
Yun Soo Myung, Yong-Wan Kim, Young-Jai Park,
Thermodynamics and phase transitions in the Born-Infeld-anti-de Sitter black holes,
Phys. Rev. D 78(2008)084002.



\bibitem{sh}
Ahmad Sheykhi, Doa Hashemi Asl, Amin Dehyadegari,
Conductivity of higher dimensional holographic superconductors with nonlinear electrodynamics,
Phys. Lett. B 781(2018)139.






\bibitem{Rogatko1}
Bart{\l}omiej Kiczek, Marek Rogatko, Karol I. Wysokinski,
Holographic DC SQUID in the presence of dark matter, JCAP 01(2021)063.



\bibitem{mr1}
D. N$\acute{u}\tilde{n}$ez, H. Quevedo, and D. Sudarsky,Black Holes Have No Short Hair, Phys. Rev. Lett.
76, 571 (1996).



\bibitem{mr2}
S. Hod,Hairy Black Holes and Null Circular Geodesics,Phys. Rev. D 84, 124030 (2011).


\bibitem{fc}
Yan Peng, Upper bound on the radii of regular ultra-compact star photonspheres,
Phys. Lett. B 790(2019)396-399.



\bibitem{Basu}
Pallab Basu, Pankaj Chaturvedi, Prasant Samantray,
Chaotic dynamics of strings in charged black hole backgrounds, Phys. Rev. D 95(2017)066014.



\bibitem{Rogatko}
Bart{\l}omiej Kiczek, Marek Rogatko,Ultra-compact spherically symmetric dark matter charged star objects,
JCAP 09(2019)049.




\bibitem{Peng Wang}
Peng Wang, Houwen Wu, Haitang Yang,
Thermodynamics and Phase Transition of a Nonlinear Electrodynamics Black Hole in a Cavity,
JHEP 07(2019)002.





\bibitem{mr5}
Arfken, G. ,Series Solutions--Frobenius Method, 8.5 in Mathematical Methods
for Physicists, 3rd ed. Orlando, FL: Academic Press, pp.454-467,1985.






\end{thebibliography}

\documentclass[aps,showpacs,preprintnumbers,amsmath, amssymb]{revtex4}

\usepackage{float}
\usepackage{graphics,epsfig}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}

\begin{document}
\baselineskip=0.8 cm

\title{{\bf A no-go theorem for scalar fields with couplings from Ginzburg-Landau models}}
\author{Guohua Liu$^{1}$\footnote{liuguohua1234@163.com}}
\author{Yan Peng$^{2}$\footnote{yanpengphy@163.com}}
\affiliation{\\$^{1}$ School of Physics and Physical Engineering, Qufu Normal University, Qufu, Shandong, 273165, China}
\affiliation{\\$^{2}$ School of Mathematical Sciences, Qufu Normal University, Qufu, Shandong 273165, China}

\vspace*{0.2cm}
\begin{abstract}
\baselineskip=0.6 cm
\begin{center}
{\bf Abstract}
\end{center}

Recently Hod proved a no-go theorem that
static scalar fields cannot form
spherically symmetric boson stars
in the asymptotically flat background.
On the other side, scalar fields can be coupled to the gradient
according to next-to-leading order Ginzburg-Landau models.
In the present work, we extend Hod's discussions by considering
couplings between static scalar fields and the field
gradient. For a non-negative coupling parameter,
we show that there is no asymptotically flat spherically symmetric
boson stars made of coupled static scalar fields.

\end{abstract}

\pacs{11.25.Tq, 04.70.Bw, 74.20.-z}\maketitle
\newpage
\vspace*{0.2cm}

\section{Introduction}

There is growing evidence supporting the existence
of fundamental scalar fields in nature. The systems composed
of scalar fields have attracted a lot of attention
from mathematicians and physicists. One of the most famous
properties of such systems is the no hair theorem,
which states that static scalar fields cannot exist
outside asymptotically flat black hole horizons,
see references \cite{Bekenstein}-\cite{Brihaye} and reviews \cite{Bekenstein-1,CAR}.
Lately, it was found that no scalar field property
also appears in curved horizonless spacetimes \cite{Hod-6}-\cite{LR1}.

A well known horizonless configuration is the boson star made of
scalar fields. Theoretically, the scalar fields can be stationary or static.
It was found that boson stars
can be constructed with stationary scalar fields \cite{FE,DAER}.
Then it is interesting to examine whether
static scalar fields can form boson stars.
In the flat spacetime, boson stars cannot be made of
static scalar fields due to Derrick's theorem \cite{GHD}.
In the asymptotically flat background, Hod proved that
boson stars cannot be composed of static scalar fields \cite{ng1,ng2}.
In the asymptotically AdS background,
this no boson star property also appears \cite{ng3}.
On the other side, in next-to-leading order Ginzburg-Landau
family of models for a BCS superconductor, there is
a new term with scalar fields coupled to the gradient \cite{cond1,cond2,cond3}
and the couplings play an important role in scalar condensations \cite{GLYP}.
Along this line, it is interesting to study the no
boson star property with coupled static scalar fields.

In the following, we plan to extend the discussion
in \cite{ng1} by considering static scalar fields
coupled to the gradient in the asymptotically flat background.
We show that boson stars cannot be made of static scalar fields
with a non-negative coupling parameter.
We summarize main results in the last section.

\section{No static boson star for coupled scalar fields}

In the extended Ginzburg-Landau models, a new term with scalar fields
coupled to the gradient appears \cite{cond1,cond2,cond3}.
In this work, we study the system with coupled static scalar fields
in the curved spacetime. The Lagrangian density describing
scalar fields coupled to the gradient is \cite{dg,sh,Rogatko1}
\begin{eqnarray}\label{lagrange-1}
\mathcal{L}=R-|\nabla_{\alpha} \psi|^{2}-\xi\psi^{2}|\nabla_{\alpha} \psi|^{2}-V(\psi^{2}),
\end{eqnarray}
where R corresponds to the Ricci curvature and $\xi$ is the coupling parameter.
In order to obtain results in this work, we assume that the coupling parameter
is non-negative as $\xi>0$. We study radial direction depending scalar
fields expressed as $\psi=\psi(r)$.

The potential $V(\psi^2)$ is positive semidefinite
and increases as a function of $\psi^2$ satisfying
\begin{eqnarray}\label{lagrange-1}
V(0)=0~~~~and~~~~\dot{V}=\frac{dV(\psi^2)}{d(\psi^2)}> 0.
\end{eqnarray}
The free scalar fields with $V(\psi^2)=\mu^2\psi^2$
apparently satisfy the relation (2) as
$V(0)=0$ and $\dot{V}=\mu^2>0$. Here $\mu$
is the nonzero scalar field mass.

The metric of spherically symmetric boson star can be expressed as
 \cite{mr1,mr2,fc,Basu,Rogatko,Peng Wang}
\begin{eqnarray}\label{AdSBH}
ds^{2}&=&-ge^{-\chi}dt^{2}+\frac{dr^{2}}{g}+r^{2}(d\theta^2+sin^{2}\theta d\phi^{2}).
\end{eqnarray}
$\chi$ and $g$ are functions of the coordinate r.
$\theta$ and $\phi$ are angular coordinates.

The Lagrangian density (1) yields the scalar field equation
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi=0.
\end{eqnarray}

At spatial infinity, metric functions asymptotically behave as \cite{ng1}
\begin{eqnarray}\label{AdSBH}
\chi\rightarrow 0,~~~~~~g\rightarrow 1~~~~~~for~~~~~~r\rightarrow \infty.
\end{eqnarray}

And near the origin, metric functions are characterized by \cite{ng1}
\begin{eqnarray}\label{AdSBH}
\chi'\rightarrow 0,~~~~~~g\rightarrow 1+O(r^2)~~~~~~for~~~~~~r\rightarrow 0.
\end{eqnarray}

The corresponding energy density is
\begin{eqnarray}\label{BHg}
\rho=-T^{t}_{t}=g\psi'^{2}+\xi g\psi^{2}\psi'^{2}+V(\psi^{2}).
\end{eqnarray}

The finite gravitational mass condition $M=\int_{0}^{\infty}4\pi r^{2}\rho dr< \infty$
implies that $r^{2}\rho$ decreases faster than $\frac{1}{r}$ around the infinity.
Then we arrive at the asymptotical behavior
\begin{eqnarray}\label{BHg}
r^{3}\rho\rightarrow 0~~~~~for~~~~~r\rightarrow \infty.
\end{eqnarray}

According to relations (2), (7) and (8), we obtain the
infinity vanishing condition
\begin{eqnarray}\label{InfBH}
&&\psi(\infty)=0.
\end{eqnarray}

The scalar field equation near the origin is
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+\frac{2}{r}(1+\xi\psi^2)\psi'-(\xi \psi'^{2}+\dot{V})\psi=0.
\end{eqnarray}

Around the origin, we set the scalar field in the general form $\psi(r)=r^{s}(a_{0}+a_{1}r+a_{2}r^2+\ldots)$ \cite{mr5}.
Putting this expression into equation (10) and considering the leading order, we find
the relation $s=0$ and the solution satisfies the near origin asymptotical behavior
\begin{eqnarray}\label{BHg}
\psi(r)=a[1+\frac{\dot{V}(a^2)}{6(1+\xi a^2)}\cdot r^2]+O(r^3),
\end{eqnarray}
where $a$ is the scalar field value $\psi(0)$ at the origin.

In the case of $a=\psi(0)=0$, also considering the relation (9),
we deduce that the eigenfunction $\psi(r)$ must possess at least
one extremum point $r_{peak}$ \cite{Hod-6}.
The scalar fields are characterized by
\begin{eqnarray}\label{InfBH}
\{ \psi^2>0,~~~~\psi\psi'=0~~~~and~~~~\psi\psi''\leqslant0\}~~~~for~~~~r=r_{peak}.
\end{eqnarray}
At $r=r_{peak}$, the relation (12) yields an inequality
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi^2<0.
\end{eqnarray}

In the case of $a>0$, (11) yields $\psi'(0)=0$ and $\psi''(0)=\frac{a\dot{V}(a^2)}{3(1+\xi a^2)}>0$,
which implies $\psi' > 0$ for $r>0$ in the near origin region.
The scalar field eigenfunction $\psi(r)$ increases to be more positive around
the origin and finally approaches zero at the infinity.
So one deduces that the scalar field $\psi(r)$ reaches
a positive local maximum value at one extremum point $\tilde{r}_{peak}$.
At this point, the eigenfunction is characterized by
 \begin{eqnarray}\label{InfBH}
\{ \psi>0,~~~~\psi'=0~~~~and~~~~\psi''\leqslant0\}~~~~for~~~~r=\tilde{r}_{peak}.
\end{eqnarray}
At $r=\tilde{r}_{peak}$, the relation (14) yields the inequality
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi<0.
\end{eqnarray}

And in another cases with $a<0$, one can deduce the
existence of one extremum point $\bar{r}_{peak}$,
where the scalar field $\psi(r)$ reaches
a negative local minimum value.
At this point, the relation is
\begin{eqnarray}\label{InfBH}
\{ \psi<0,~~~~\psi'=0~~~~and~~~~\psi''\geqslant0\}~~~~for~~~~r=\bar{r}_{peak}.
\end{eqnarray}
At $r=\bar{r}_{peak}$, the characteristic inequality is
\begin{eqnarray}\label{BHg}
(1+\xi\psi^2)\psi''+[(1+\xi\psi^2)(\frac{2}{r}-\frac{\chi'}{2}+\frac{g'}{g})+2\xi \psi \psi']\psi'-(\xi \psi'^{2}+\frac{\dot{V}}{g})\psi>0.
\end{eqnarray}

Relations (13), (15) and (17) are in contradiction with the scalar field equation (4).
So the scalar field equation cannot be respected at the corresponding extremum
points. We therefore conclude that
spherically symmetric asymptotically flat boson stars cannot
be constructed with static coupled scalar fields with
non-negative coupling parameters.

\section{Conclusions}

We studied the nonexistence of boson stars composed
of static scalar fields in the spherically symmetric
asymptotically flat background. We considered static
scalar fields coupled to the field gradient,
where the coupling term also appears in the
next-to-leading order Ginzburg-Landau models. For non-negative
coupling parameters, we obtained the characteristic
inequalities (13), (15) and (17) at corresponding extremum points.
However, these relations are in contradiction
with the static scalar field equation (4).
As a summary, we proved the nonexistence of
spherically symmetric asymptotically flat
static boson stars made of scalar fields with
a non-negative coupling parameter.

\begin{acknowledgments}

This work was supported by a grant from Qufu Normal University
of China under Grant No. xkjjc201906. This work was also supported by the Youth Innovations and Talents Project of Shandong
Provincial Colleges and Universities (Grant no. 201909118).

\end{acknowledgments}

\end{document}